# and quantum information science

### v1

@@ 5.3 Constraints on the cost function from expected time dynamics
In Sec. 4.1, we discretized time to show that quantum complexity exhibits two behaviors: linear growth for an exponential long time and the switchback effect. These derivations relied on simple counting arguments. Afterward, we quantitatively showed that complexity exhibits linear growth, using random unitary circuits, algebraic geometry, and unitary designs (Sections 4.34.6). We aim to reproduce the above key features of complexity, using complexity geometry. In other words, we ask the question: what are necessary conditions for a complexity geometry to reproduce linear growth and the switchback effect when applied to the unitary unitary time evolution generated by a chaotic Hamiltonian? To do so, we select an appropriate metric on the unitary manifold $\mathrm{SU}(N)$ (assuming $N\gg 1$). The associated cost function should satisfy the following requirements [80, 39, 44]:
@@
    The group manifold has a negative average-sectional-curvature [80, 39], such that nearby geodesics diverge. This property is necessary (but insufficient) for reproducing chaotic and ergodic behaviors of the Hamiltonians that generate motion on the unitary manifold.
@@
    The Nielsen complexity displays the switchback effect, which requires the geometry’s sectional curvatures to scale as $1/K$ ($K$ denotes the number of qubits).
@@
Let us comment on the implications of these requirements. First, consider the minimization, in Eq. (23), over the stabilizer group of $\lvert\psi(t)\rangle$, yielding a metric on the space of states rather than on unitaries. Do the above properties reflect in this geometry on the space of states? The minimization breaks the geometry’s homogeneity, complicating one’s study of the geodesics and curvature in the space of states. [Granted, the minimization projects the metric onto a manifold of lower dimensionality than SU$(N)$’s.] Nonetheless, as mentioned earlier, for the case of Riemannian cost functions, the map (22) is a Riemannian submersion [86] (defined in [87]). Riemannian submersions interrelates the sectional curvatures in the space of unitaries and the sectional curvatures in the space of states, via the O’Neill formula [100]. The O’Neill formula implies that the sectional curvature along a plane in the space of states is lower-bounded by the sectional curvature along an appropriate plane in the unitary manifold.
@@
Second, the group manifold equipped with the Cartan-Killing metric—the standard metric on SU$(N)$—does not satisfy the above requirements because it has only constant positive sectional curvatures. How can one construct a geometry that has negative sectional curvatures? One can assign the penalty factors such that the $\mathfrak{su}(N)$ algebra’s nonvanishing commutators have the structure $[easy,\,easy]=\,hard$. The easy and hard refer to generators associated with small and large penalty factors in the cost function [37]. When the penalty factors are associated to each generator such that the commutators take the above-mentioned form, then sectional curvatures’ negativity follows from the Pythagorean theorem for curved space. The latter statement was proven in the contexts of a qubit’s unitary manifold and a qutrit’s [37, 86]. [^13] A qudit is a $d$-level quantum system, represented by a Hilbert space $\mathbb{C}^{d}$. A qubit has $d=2$; and a qutrit, $d=3$. Manifolds equipped with right-invariant metrics must have positive sectional curvatures in some regions; otherwise, they are flat [101]. The geodesic’s length computed in flat space does not present the features expected from a measure of complexity, therefore it is not interesting for our purposes.
@@
References [96, 98] studied time-dependent trajectories generated by the SYK Hamiltonian on the unitary manifold. The authors showed that the unitary’s geodesic distance from the origin is lower-bounded by a geodesic whose length grows linearly with time. This linear growth is truncated at conjugate points. At a conjugate point, the geodesic does not result from extending an earlier-time geodesic; a shorter path connects the origin to the unitary. Conjugate points in the critical complexity geometry described at the end of Sec. 5.2 are pushed to infinity [82]. Therefore, the linear growth of Nielsen’s complexity in that geometry persists for a time exponential in the number of qubits. In the context of spin chains, Brown showed that many complexity geometries’ diameters are exponentially large in the number of qubits [14]. This exponential relation is necessary for the complexity to saturate at an exponentially large value (Fig. 3).
@@
We have argued that Nielsen’s complexity exhibits one behavior characteristic of quantum complexity: linear growth. Nielsen’s complexity reproduces the other behavior—the switchback effect—in toy models [44]. Qualitative features similar to the switchback effect are also displayed by single-qubit and two-qubit systems [102]. However, the comprehensive picture of the emergence of the switchback effect from the complexity geometry is still lacking and is one of the important open questions that future research in this area should address.

### v2

@@ 5.3 Constraints on the cost function from expected time dynamics
In Sec. 4.1, we showed that quantum complexity for discrete time circuits exhibits two behaviors: linear growth which can last for an exponentially long time, and the switchback effect. These derivations relied on simple counting arguments. Afterwards, we quantitatively demonstrated this linear growth for random unitary circuits, by using algebraic geometry and unitary designs (Sections 4.34.6). We would like to understand which complexity geometries reproduce the above key features of complexity when applied to the unitary time evolution generated by a chaotic Hamiltonian. To do so, we must select an appropriate metric on the $\mathrm{SU}(N)$ manifold. The requirements turn out to be [95, 50, 55]:
@@
    The group manifold must have a negative average sectional curvature [95, 50], such that nearby geodesics diverge. This is necessary, but not sufficient, for establishing the above key features for chaotic and ergodic Hamiltonians.
@@
    The geometry’s sectional curvatures should scale as $1/K$, where $K$ denotes the number of qubits, in order for the Nielsen complexity to display the switchback effect.
@@
Let us comment on the implications of these requirements. First, consider Eq. (23) which yields a metric on the space of states rather than on unitaries. Do the above properties have a reflection in the induced geometry on the space of states? In fact the induced metric is not homogeneous, complicating the study of the geodesics and curvature in the space of states [108]. Nonetheless, as mentioned earlier, for the case of Riemannian cost functions, the map (22) is a Riemannian submersion [101] (defined in [102]). Riemannian submersions relate sectional curvatures in the space of unitaries and sectional curvatures in the space of states, via the O’Neill formula [116]. The latter implies that the sectional curvature along a plane in the space of states is lower-bounded by the sectional curvature along an appropriate plane in the unitary manifold.
@@
Second, the group manifold equipped with the Cartan-Killing metric—the standard metric on SU$(N)$—does not satisfy the above requirements because it has only constant positive sectional curvatures. How can we construct a geometry that has negative sectional curvatures? We can assign the penalty factors such that some of the $\mathfrak{su}(N)$ algebra’s nonvanishing commutators have the structure $[easy,\,easy]=\,hard$. The easy and hard refer to generators associated with small and large penalty factors in the cost function. When the penalty factors are associated to generators such that the commutators take the above-mentioned form, then the Pythagorean theorem for curved space guarantees some negative sectional curvatures [108]. The latter statement was proven in the contexts of the unitary manifolds of qubits and qutrits [108, 101]. [^12] A qudit is a $d$-level quantum system, represented by a Hilbert space $\mathbb{C}^{d}$. A qubit has $d=2$; and a qutrit, $d=3$. Note that manifolds that are equipped with right-invariant metrics, like the ones we are interested in, must also have some positive sectional curvatures; otherwise, they are flat [117].
@@
References [112, 114] studied time-dependent trajectories generated by the SYK Hamiltonian on the unitary manifold. The authors showed that the geodesic distance from the origin to the time evolution operator is lower-bounded by a geodesic whose length grows linearly with time. This linear growth is truncated at conjugate points along the geodesic. There is a shorter path connecting the origin to a conjugate point than the one produced by extending the earlier time geodesic. It is possible that the linear growth is truncated even earlier by global obstructions – geodesics that wind in some other direction around the unitary group to the same point – but these are much more difficult to study, as we will discuss in Sec. 5.4. Conjugate points in the critical complexity geometry described at the end of Sec. 5.2 are pushed to infinity [97]. Therefore, the linear growth of Nielsen’s complexity in that geometry persists for a time exponential in the number of qubits. In the context of spin chains, Brown showed that many complexity geometries have diameters that are exponentially large in the number of qubits [19]. This exponential relation is necessary for complexity to saturate at an exponentially large value (Fig. 3).
@@
We have argued that Nielsen complexity exhibits one behavior characteristic of quantum complexity: the initial linear growth. Nielsen complexity also reproduces the other behavior—the switchback effect—in toy models [55]. Qualitative features similar to the switchback effect are also displayed by single-qubit and two-qubit systems [118]. However, a comprehensive picture of the emergence of the switchback effect from the complexity geometry is still lacking and is an open question for future research.
5.4 Physical criteria for complexity growth: general results and applications
How is Nielsen complexity related to standard concepts in physics, like energy spectra, correlation functions, and Schrödinger dynamics? Several connections were uncovered by the authors of [112, 114], who investigated physical criteria that control the length of time over which Nielsen complexity will grow during Hamiltonian time evolution. We will discuss their results below, and use this discussion as a way of illustrating how to perform practical calculations of Nielsen complexity.
For concreteness, let us consider a time-independent Hamiltonian acting on an $N$–dimensional Hilbert space. Time evolution is then governed by the unitary operator $U(t)=e^{-iHt}$. We want to think of this operator as tracing a trajectory on the $\mathrm{SU}(N)$ group manifold from the identity $\mathbbm{1}$ at $t=0$ to $U(t)$. Following Nielsen, we will regard the complexity of $U(t)$ as measured by the length of the shortest geodesic from $\mathbbm{1}$ to $U(t)$ in a certain complexity metric that we will define below.
In fact, finding the shortest geodesic on $\mathrm{SU}(N)$ for general $N$ is a difficult problem in any metric because the group manifold has a complicated fiber bundle topology. So, if we find one geodesic between two group elements by solving the appropriate differential equation, there may well be others that start at the identity and go “around” the group in another direction to get to the same point. Some of the latter may even be shorter. The simplest example of this occurs in $\mathrm{SU}(2)$, whose manifold is a 3-sphere. Geodesics on the 3-sphere are great circles. If we continue on a great circle past the point diametrically opposite to the initial one, there is a shorter geodesic that goes the other way around the sphere. For general $\mathrm{SU}(N)$ there may be many more distinct paths between two group elements that exploit the complicated topology, and finding them all is an unsolved problem. The problem is even believed to be NP-complete from an algorithmic perspective. While finding all global shortcuts on $\mathrm{SU}(N)$ equipped with a given complexity metric is intractable, the authors of [112, 114] pointed out that there is a class of shortcuts that we can analyze, associated to conjugate points along a geodesic. Conjugate points imply that the geodesic we started with is a saddlepoint, not a minimum of the length, so that we can find a shorter path by descending from the saddle. Below we will explain how to find conjugate points, and present physical criteria that delay their occurrence until exponential times under Hamiltonian evolution.
Our goal is to find geodesics on a Lie group manifold in a metric that penalizes “hard” generators that represent infinitesimal transformations that are difficult to implement. First, consider the group generators that span the tangent space to the group manifold at the identity. Consider the Cartan-Killing form $K_{IJ}\propto f_{IM}^{L}f_{JL}^{M}$ where $f_{IJ}^{L}$ are the group’s structure constants defined using the commutation relation for the group generators $[T_{I},T_{J}]=if_{IJ}^{L}T_{L}$ with repeated indices summed. Now define
$G_{IJ}=\frac{c_{I}+c_{J}}{2}K_{IJ}\,,$ (26)
with $c_{I}=\mathcal{O}(1)$ for “easy” directions in the group’s tangent space at the identity, and $c_{I}=\mathcal{O}(e^{\epsilon S})$ for the “hard” directions. Here $\epsilon$ is any constant, and $S=\ln N$ is the logarithm of the dimension of the Hilbert space on which the group acts. We may take the hard directions to be generators that act on more than $k$ qubits for some $k$, or act non-locally, or are differentiated in some other way appropriate to the problem at hand. For $\mathrm{SU}(N)$ we will split generators into the easy ones $\{T_{\alpha}\}$ and the hard ones $\{T_{\dot{\alpha}}\}$, so that the diagonal metric is
$G_{IJ}=\begin{pmatrix}\delta_{\alpha\beta}&0\\ 0&(1+\mu)\,\delta_{\dot{\alpha}\dot{\beta}}\end{pmatrix}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mu\sim e^{\epsilon S}\,.$ (27)
This is the metric, or inner product, on the tangent space at the identity. If $X$ and $Y$ are tangents at any other point $U$ on the group manifold we define $g(X,Y)=G(XU^{-1},YU^{-1})$ where $G$ is a functional notation for the bilinear inner product defined in (26).
With these definitions, the geodesic equation on $\mathrm{SU}(N)$ in Euler-Arnold form is [112] [^13] We relabeled the time $t$ along the circuit as $s$ to avoid confusion with the physical time of the system.
$G_{IJ}\frac{dV^{J}}{ds}=f_{IJ}^{K}\,V^{J}G_{KL}V^{L}\,,$ (28)
with repeated indices summed. Here, $V^{I}$ is tangent to the path, [^14] The $V^{I}$ are analogous to the control functions $Y^{I}$ in equation (17); we use $V^{I}$ in this subsection and reserve $Y$ for the map described below Eq. (34). $s$ is the parameter along the path, and $f_{IJ}^{K}$ are the structure constants of the group. If we solve this for the tangents $V^{I}(s)$, then the geodesic path is
[…]
where $\mathcal{\overset{\leftarrow}{P}}$ represents path ordering. We consider the Nielsen complexity defined by
[…]
where the minimum is taken over all geodesics from $I$ to $U(s)$. This is a generalization of the complexity definition associated with the $p=2$ cost function in Eq. (25).
This is a practical formalism for calculations. For example, consider the simple case of $\mathrm{SU}(2)$ with generators $T_{1,2}=\gamma_{1,2}$ and $T_{3}=i\gamma_{1}\gamma_{2}$, where $\gamma_{i}$ are the standard Gamma matrices and $T_{3}$ is deemed “hard”, while $T_{1,2}$ are easy [112]. Geodesics take the explicit form
$V^{1}(s)=v^{1}\cos(v^{3}\mu s)-v^{2}\sin(v^{3}\mu s)\leavevmode\nobreak\ \leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\ V^{2}(s)=v^{2}\cos(v^{3}\mu s)+v^{1}\sin(v^{3}\mu s)\leavevmode\nobreak\ \leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\ V^{3}(s)=\frac{v^{3}}{2}\,,$ (31)
and the Nielsen complexity of going from $U(1)=\mathbbm{1}$ to $U(1)=U_{{\rm target}}$ is
$\mathcal{C}=\sqrt{(v^{1})^{2}+(v^{2})^{2}+(1+\mu)(v^{3})^{2}}\,,$ (32)
where the initial velocities $v^{I}$ have to be chosen to get to the target at $s=1$. We can now pick geodesics for which $U(1)=e^{-iHt}$ with a Hamiltonian $H=J_{1}T_{1}+J_{2}T_{2}$ built from the “easy” generators. It is straightforward to then show that the complexity initially grows linearly as $\mathcal{C}(t)=t\sqrt{J_{1}^{2}+J_{2}^{2}}$ as the Hamiltonian itself generates the shortest trajectory from the identity by avoiding the more expensive direction $T_{3}$. The linear growth ends when the trajectory reaches the opposite pole of $\mathrm{SU}(2)$ from the identity because then there is a shorter geodesic to $e^{-iHt}$ that goes around the other side of the $S^{3}$ manifold of $\mathrm{SU}(2)$. This causes the complexity to decrease back to zero as $U(t)=e^{-iHt}$ returns to the origin. Continued time evolution leads to oscillations in the complexity. We can view the return to zero complexity as the analog of a recurrence in this simple integrable dynamics. Notice that the linear growth of complexity was truncated by a shortcut in the unitary group. In this case the shortcut is a global obstruction to complexity growth – it exists because of the non-trivial global topology of the group manifold.
Now consider $\mathrm{SU}(N)$ for general $N$. Following [112, 114] we will consider a representation of the group based on the algebra of Gamma matrices $\gamma_{a}$ with $a=0,\dots,N-1$ and $\{\gamma_{a},\gamma_{b}\}=2\delta_{ab}$. We can construct the $2^{N}$ generators of $\mathrm{SU}(N)$ as $T_{a_{1}\cdots a_{m}}=\gamma_{a_{1}}\cdots\gamma_{a_{m}}$ subject to the condition $a_{p}<a_{q}$ if $p<q$. We will label the generators as $T_{I}$ with a multi-index $I\equiv(a_{1}\cdots a_{m})$. We will later discuss a physical realization of the SYK model, constructed from $N$ Majorana fermions $\psi_{a}$ where $\gamma_{a}\sim\psi_{a}$. In view of this identification, we will say that generators $T_{\alpha}$ with no more than $k$ $\gamma$ matrices in their definition are k-local, and that generators $T_{\dot{\alpha}}$ with more than $k$ $\gamma$ matrices in their definition are k-nonlocal. We will also say that k-local generators are “easy” and that k-nonlocal generators are “hard” and are therefore penalized by $1+\mu$ in the metric in (27). Now suppose that the Hamiltonian is k-local: $H=\sum_{\alpha}J^{\alpha}T_{\alpha}$. Then the authors of [112, 114] showed from the Euler-Arnold equation (28) that there is always a geodesic from the identity to $e^{-iHt}$ along the path taken by the physical time evolution. In the notation of (28), the components of the tangent vectors along this geodesic are $V^{\alpha}=J^{\alpha}t$ leading to the tangent vector $V=Ht$. This means that the Nielsen complexity grows linearly in time,
$\mathcal{C}(t)=t\,\sqrt{\sum_{\alpha}(J^{\alpha})^{2}}$ (33)
if the time evolution trajectory, which is a geodesic, is also the shortest geodesic.
At least at the early stages of time evolution, we can be sure from the results discussed above that the trajectory generated by the Hamiltonian is the shortest between $\mathbbm{1}$ and $e^{-iHt}$. How could this linear geodesic trajectory stop being the shortest? One possibility is the existence of a geodesic loop that allows us to exploit the topology of $\mathrm{SU}(N)$ to arrive at the same destination through a different path. This is what happened in the $\mathrm{SU}(2)$ example above. Geodesic loops are difficult to work out in the general unitary group, so we will not attempt to do that. However, note that in $\mathrm{SU}(2)$ the shorter path arising from a great cycle on the group manifold became relevant after the time evolution trajectory had traversed a diameter of the manifold. We expect something like that to be also generally true for geodesic loops of $\mathrm{SU}(N>2)$. Following Ref. [19], we also expect the diameter of $\mathrm{SU}(N)$ to be exponentially large in the number of degrees of freedom in complexity metrics satisfying reasonable consistency criteria. This means that, all else being equal, we expect the geodesic loop obstruction to start to play a role in complexity growth after exponential time. So we will focus on another kind of shortcut on the group manifold that could truncate the linear complexity growth in (33) far earlier. These shortcuts are associated to conjugate points on the linear Hamiltonian trajectory that we will now describe.
A conjugate point occurs on a geodesic $U(s):[0,1]\to\mathrm{SU}(N)$ with $U(0)=P$ and $U(1)=Q$ if a perturbation of the path produces a new curve with the same boundary conditions that satisfies the geodesic equation to first order. This implies that the geodesic we started with is a saddlepoint, not a minimum of the length, so that we can find a shorter path. The actual minimizing geodesic signaled by the appearance of conjugate point can be a finite distance away. Since conjugate points indicate converging geodesics, they cannot occur if all the sectional curvatures are negative. There is a result of Milnor [119] that says that most sectional curvatures of Lie groups are negative. So some authors have considered high genus Riemann surfaces with a metric induced from the covering space, a hyperbolic disk, as a toy model of the unitary group manifold [50]. In such a toy model there are no conjugate points, so they are irrelevant. Others have constructed special metrics on the unitary group that are designed to push the conjugate points to infinity [97]. But these approaches seem misleading about the structure of the unitary group because the main result of Milnor is that any right invariant metric on $\mathrm{SU}(N)$ with $N>2$ must have some positive sectional curvature, or else it is flat. Indeed, as we will see, conjugate points are some of the earliest obstructions to complexity growth in free and integrable theories.
To find conjugate points we start with a linear path along the time evolution trajectory from $\mathbbm{1}$ to $e^{-iHt}$. This path has tangents $V(s)=Hs$ which produce a line of unitaries $U(s)=e^{-istH}$ with $s=[0,1]$. We want to perturb the trajectory $V(s)\to V(s)+\delta V(s)$ and still solve the Euler-Arnold equation to first order, with the same boundary conditions. In first order perturbation theory, this gives the Jacobi equation. In detail, let $\delta V=\delta V_{L}+\delta V_{NL}$ where the subscript $L$ indicates a projection to the k-local/easy subspace, and the subscript $NL$ denote a projection to the k-nonlocal/hard subspace. Then the Jacobi equation implies:
$i\frac{d\delta V_{L}(s)}{ds}=\mu t[H,\delta V_{NL}(s)]_{L}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ i\frac{d\delta V_{NL}(s)}{ds}=\frac{\mu t}{1+\mu}+[H,\delta V_{NL}(s)]_{NL}$ (34)
where the subscripts $L$ and $NL$ again indicate projection to the k-local/easy and k-nonlocal/hard subspaces.
By solving this equation, we can work out how an initial perturbation of the trajectory at the origin changes the trajectory. The authors of [112] used the formal solution of the above equation to write down a superoperator $Y[\delta V(0)]$ mapping the initial perturbation of the trajectory, $\delta V(0)$, onto a perturbation of the endpoint. Conjugate points occur when the endpoint is unchanged, namely when the superoperator $Y$ has a zero mode in the formalism of [112]. We do not write the superoperator explicitly here because it has a complicated form that requires further definitions to explain, and instead refer the reader to [112]. By working out criteria for the appearance of such zero modes, the authors of [112, 114] arrived at a number of general results for the appearance of conjugate points that are described below:
  1. Result 1:
    For arbitrary local Hamiltonians with a finite complexity cost factor $\mu$, conjugate points must exist at finite distance along the linear geodesic. This is proved by starting with the result of [120] that conjugate points exist at finite distance when the complexity penalty $\mu$ vanishes. The persistence of the conjugate points is then established by a continuity argument as $\mu$ increases, along with Morse theory on the space of paths on the group [114]. So the absence of conjugate points must be shown to establish linear complexity growth until times of order $t\sim e^{S}$.
  2. Result 2:
    If a q-local Hamiltonian has an adjoint eigenoperator $O$ such that $[H,O]=\lambda O$ with $\lambda\in\mathbb{R}$ and $O\in\{{\rm local/easy\ operators}\}$, then conjugate points occur at times $t_{*}=\frac{2\pi}{\lambda}\mathbb{Z}$ (where $\mathbb{R}$ are the reals and $\mathbb{Z}$ are the integers). This is proven in [114] by explicit examination of the zero mode condition on the superoperator $Y$ mentioned above.
  3. Result 3:
    If a q-local Hamiltonian has an adjoint eigenoperator $O^{\prime}$ such that $[H,O^{\prime}]=\lambda^{\prime}O^{\prime}$ with $\lambda^{\prime}\in\mathbb{R}$ and $O^{\prime}\in\{{\rm nonlocal/hard\ operators}\}$, then conjugate points occur at times $t_{*}=\frac{2\pi(1+\mu)}{\lambda^{\prime}}\mathbb{Z}$. This is proven in the same way as Result 2 above.
  4. Result 4:
[…]
  5. Result 5:
    Let $\lvert m\rangle$ and $\lvert n\rangle$ be any two energy eigenstates, and let $T_{\alpha}$ be the local/easy generators, and $T_{\dot{\alpha}}$ the nonlocal/hard generators. Then linear growth of Nielsen complexity persists for a time of $\mathcal{O}(e^{\epsilon S})$, where $S$ is the logarithm of the Hilbert space dimension, if
    $R_{mn}=\frac{\sum_{\alpha}|\langle m\rvert T_{\alpha}\lvert n\rangle|^{2}}{\sum_{\alpha}|\langle m\rvert T_{\alpha}\lvert n\rangle|^{2}+\sum_{\dot{\alpha}}|\langle m\rvert T_{\dot{\alpha}}\lvert n\rangle|^{2}}=e^{-2S}{\rm poly}(S)r_{mn}$ (35)
    where $r_{mn}=\mathcal{O}(1)$ for $m\neq n$ and ${\rm poly}(S)$ means polynomial in $S$. This criterion, which is called the Eigenstate Complexity Hypothesis (ECH) (since it recalls the Eigenstate Thermalization Hypothesis), says that energy eigenstates cannot be mapped onto each other by easy operators. This result was proven in [112] by showing that if we assume ECH, the zero modes of the superoperator $Y$ occur at times $t^{*}=(2\pi/\Delta_{{\rm max}})(1+e^{\epsilon S})$, where $\Delta_{\rm max}$ is the largest gap in the spectrum.
These theorems can be applied to the SYK model with $n$ Majorana fermions, which we will return to in Sections 6.3.3 and 7.5.1. Here we briefly summarize the results of [112, 114] in this regard. The authors provide evidence that: (a) The Nielsen complexity of the free SYK model (with only quadratic interactions) is bounded as $\mathcal{C}(t)\lesssim\sqrt{n}$; (b) The Nielsen complexity of integrable deformations of the free SYK model is bounded as $\mathcal{C}(t)\lesssim{\rm poly}(n)$, meaning that the initial linear growth is guaranteed to truncate in polynomial time; (c) 35 above, and numerical analysis of the spectrum and eigenstates of chaotic deformations of the free SYK model suggest that the Nielsen complexity of such systems is bounded as $\mathcal{C}(t)\lesssim e^{\epsilon n}$, meaning that linear complexity growth can persist for exponential time, if we neglect possible global obstructions from geodesic loops. The finding (c) applies to any theory in which the energy eigenstates are a Haar random unitary rotation of a “standard” or “free” basis of states. The ETH hypothesis for chaotic theories suggests that their eigenstates may indeed be of this form.